OpenStudy (shelby1290):

Find the value of k such that the quadratic equation kx^2 + (k+8)x+9=0 has one root

OpenStudy (dayakar):

do u have any idea about the solution

OpenStudy (shelby1290):

@dayakar I think I have to find the zeros but I just don't know how to

OpenStudy (shelby1290):

How do I solve this question?

OpenStudy (dayakar):

is it first term k^2 x^2

OpenStudy (shelby1290):

no, the first term is kx^2

OpenStudy (dayakar):

can u compare the equation with ax^2+bx+c=0

OpenStudy (dayakar):

a=k, b=(k+8)^2 , c= 9

OpenStudy (dayakar):

here one root means the equation has equal roots

OpenStudy (dayakar):

that means , discriminant =b^2-4ac =0

OpenStudy (dayakar):

do u get it

OpenStudy (shelby1290):